# Interference patterns for simple lens models in wave-optics regime

### v1

@@ 5 Chang-Refsdal lens
After the point mass lens, arguably, the next simplest lens model is a point mass in the presence of external shear caused by some external mass distribution, also known as Chang-Refsdal lens (Chang & Refsdal, 1979, 1984). The lensing potential (in dimensionless form) for such a lens model is given as,
@@
where $(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$ denote the components of the external shear. Here, we have assumed that the point mass lens is at the origin. With this, the arrival time delay surface takes the form
@@
where $M$ is the mass of the point lens, and other quantities have their usual meaning. In principle, we can always choose a coordinate system such that in this new coordinate system $\gamma_{1}^{\prime}\equiv\gamma=\sqrt{\gamma_{1}^{\prime 2}+\gamma_{2}^{\prime 2}}$ and $\gamma_{2}^{\prime}=0$. Hence, the Chang-Refsdal lens is described by a total of four free parameters, $(M_{z},\gamma,y1,y2)$. However, in the current work, we choose $\gamma_{1}^{\prime}=\gamma_{2}^{\prime}=\gamma/2$ so that caustics are somewhat diagonal in the plots for better visualization. In the absence of a point mass lens, the lensing magnification is given as $\mu_{\rm m}=1/(1-\gamma^{2})$ and based on the value of $\gamma$ the Chang-Refsdal lens represents a point mass lens sitting close to a positive parity ($\gamma<1$) or negative parity ($\gamma>1$) images formed due to the external mass distribution. Note that positive parity case ($\gamma<1$), here only corresponds to the minima image. For the maxima image, one also needs to add a constant convergence ($\kappa^{\prime}$) such that both eigenvalues of the lensing Jacobian are negative. For $\kappa^{\prime}>1$, sometimes also referred to as the over-focussing case, it is possible to have a region in the source plane where the number of lensed images is zero (e.g., Chang & Refsdal, 1984; An & Evans, 2006) and it would be interesting to see how the corresponding amplification maps look compared to other cases. In this work, for simplicity, we limit ourselves to $\gamma<1$ (i.e., minima image) and other cases are left for the future.
@@
The amplification ($|F|$) maps for three different values of $\gamma$ are shown in Figure 2 for $f=50,100,500$ Hz. We can see that, due to the breaking of circular symmetry, the oscillations in the maps are no longer circular and align with the shear direction (or major axis of the diamond caustic). For comparison, the corresponding amplification maps for $f=100$ Hz (i.e., middle column in Figure 2) under geometric optics are shown in Figure 3. Comparing Figure 2 and 3, we can see that the positions of de-amplified regions occurring with full wave-optics solution are well-traced by geometric optics approximation inside the diamond caustic as well as far from it in outer regions. As we cross the fold caustic (from inside to outside), two of the extrema points of the arrival time delay surface merge and disappear, which explains the sudden change in the $|F|$ maps under geometric optics approximation in Figure 3. On the other hand, considering the full wave-optics solution, we can see that there is still non-zero contribution from regions close to the points where the two images merge and disappear as the distortion of arrival time delay surface is continuous as a function of the source position. In addition, we can also appreciate the effect of macro-magnification $\mu_{\rm m}$. An increase in $\mu_{\rm m}$ is equivalent to increasing the point mass lens (e.g., Diego et al., 2018) bringing the oscillations in $F(f)$ at lower frequencies.
@@
Example amplification curves for source positions marked by black solid stars in the top row, black solid stars in the middle row, and green solid stars in the middle row of Figure 2 are shown in the top, middle, and bottom rows of Figure 4, respectively. For the source outside the diamond caustic and close to the cusp point in the top row of Figure 2, we have the formation of two virtual images with the minima being highly amplified, and as we saw in the point mass lens case, interference of two images with very different amplification factors although leads to frequency-dependent oscillations, however, the amplitude of these oscillations is small as seen in the blue curve in the top row of Figure 4. We also notice that solutions under geometric optics differ significantly compared to the full wave optics solution. This difference between geometric and wave optics solutions is even more prominent for the orange curve representing a source inside the diamond caustic. However, in the top panel of Figure 4, we see that the position of crests and troughs along the frequency axis matches in wave and geometric optics.
@@
In the middle row of Figure 4, we show amplification factors for source positions such that they follow a de-amplified region while crossing the diamond caustic along the major axis, whereas the bottom row represents sources which go perpendicular to the de-amplified region. The merger and disappearance of images can be seen in the $\tilde{F}(t)$ curves and the inset plots. Here, for a source inside the diamond caustic, we again see that the position of crests and troughs along the frequency axis matches in wave and geometric optics. However, once we cross the caustic, the geometric optics solution loses contribution from two images and differs significantly from the full wave optics solution.
@@ 6 Binary lens
where $q_{i}=M_{i}/M_{T}$ with $M_{T}=M_{1}+M_{2}$ represents the mass ratios of primary and secondary components and $2|\boldsymbol{L}|$ is the distance between the two lens components. Here, we have assumed that both point mass lenses are at equal distances from the origin. The corresponding arrival time delay surface is given as
@@
where various quantities have their usual meaning. Without loss of generality, we can always choose a coordinate system such that both lenses lie on the x-axis, i.e., $\boldsymbol{L}=(L,0)$. With that, we can see that a binary lens can be described by a total of five parameters, $(M_{1z},M_{2z},L,y_{1},y_{2})$. Note that the Chang-Refsdal lens is also a limiting case of binary lens (Schneider & Weiss, 1986; Dominik, 1999).
The interference patterns for a binary lens with equal mass ratio components ($M_{Tz}=200~{}{\rm M_{\odot}}$) and different $L$ values are shown in Figure 5. Example amplification factor curves for some source positions, marked by a solid star, are shown in Figure 6. Again, as we have seen in the last section, amplification factor curves can differ significantly compared to geometric optics approximation close to caustics where the number of virtual images changes (top row in Figure 6). As we move away from caustics, both the amplitude and phase of the oscillation start to converge towards geometric optics results (bottom row in Figure 6). For comparison, in the bottom panel, we also show the point mass amplification factor with $y=|\boldsymbol{y}|$. The interference patterns for a binary lens with a mass ratio  $(q_{1},q_{2})=(0.25,0.75)$ are shown in Figure 7. For binary lens, again, the dimensionless frequency $(\nu=4{\rm G}M_{Tz}f/{\rm c^{3}})$, can be used to determine other pairs of total mass and frequency (assuming the same mass ratios) which would produce the above amplification maps.
@@ 7 Summary
In this work, we simulate and look at the interference patterns for simple lens models, namely, point mass lens, Chang-Refsdal lens, and binary lens, primarily focussing on GW signals detected with the LVK detector network. Point mass lens comes with axial symmetry (and a point caustic at the centre), and the corresponding amplification factor calculated under geometric optics closely follows the full solution after the first peak (i.e., amplification region; see Bulashenko & Ubach 2022). However, once we go beyond the point mass lens, for example, by introducing external shear, the corresponding (diamond) caustics in the source plane cover a finite region of different virtual image multiplicities with the interference map showing regions of (de-)magnification aligned with the direction of shear. As the source moves close to the caustics, the corresponding amplification factor shows significant differences compared to the amplification factor obtained under geometric optics approximation due to the fact that as the source crosses a caustic, some images appear or disappear, drastically changing the contribution to geometric optics amplification factor. Caustics also mark the region around which we have multiple virtual images with similar amplification values and can lead to significant de-amplification around ${\sim}100$ Hz for a lens mass of ${\sim}100{\rm M_{\odot}}-200{\rm M_{\odot}}$. If these significantly de-amplified regions in the amplification factor align with ringdown frequencies in the LVK band, we may end up with GW signals, which may be hard to interpret without the incorporation of lensing.
@@ Acknowledgements
The author acknowledges the support of grant 2020750 from the United States-Israel Binational Science Foundation (BSF), grant 2109066 from the United States National Science Foundation (NSF), and the Ministry of Science & Technology, Israel. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.

### v2

@@ 5 Chang-Refsdal lens
After the point mass lens, the next simplest lens model is a point mass in the presence of external shear caused by some external mass distribution, also known as the Chang-Refsdal lens (Chang & Refsdal, 1979, 1984). The lensing potential (in dimensionless form) for such a lens model is given as,
@@
where $(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$ denote the components of the external shear. Here, we have assumed that the point mass lens is at the origin. With this, the arrival time delay surface takes the form,
@@
where $M$ is the mass of the point lens, and other quantities have their usual meaning. In principle, we can always choose a coordinate system such that in this new coordinate system $\gamma_{1}^{\prime}\equiv\gamma=\sqrt{\gamma_{1}^{\prime 2}+\gamma_{2}^{\prime 2}}$ and $\gamma_{2}^{\prime}=0$. Hence, the Chang-Refsdal lens is described by a total of four free parameters, $(M_{z},\gamma,y1,y2)$. However, in the current work, we choose $\gamma_{1}^{\prime}=\gamma_{2}^{\prime}=\gamma/2$ so that caustics are somewhat diagonal in the plots for better visualization. In the absence of a point mass lens, the lensing (macro-)magnification is given as $\mu_{\rm m}=1/(1-\gamma^{2})$ and based on the value of $\gamma$ the Chang-Refsdal lens represents a point mass lens sitting close to a positive parity ($\gamma<1$) or negative parity ($\gamma>1$) image formed due to the external mass distribution. Note that the positive parity case ($\gamma<1$) only corresponds to the minimum image. For the maximum image, one also needs to add a constant convergence ($\kappa^{\prime}$) such that both eigenvalues of the lensing Jacobian are negative. For $\kappa^{\prime}>1$, sometimes also referred to as the over-focussing case, it is possible to have a region in the source plane where the number of lensed images is zero (e.g., Chang & Refsdal, 1984; An & Evans, 2006) and it would be interesting to see how the amplification maps look compared to other cases. In this work, for simplicity, we limit ourselves to $\gamma<1$ (i.e., minimum image) and other cases are left for the future.
@@
The amplification ($|F|$) maps for three different values of $\gamma$ are shown in Figure 2 for $f=50,100,500$ Hz. We can see that, due to the breaking of circular symmetry, the oscillations in the maps are no longer circular and align with the shear direction (or major axis of diamond caustic). For comparison, the corresponding amplification maps for $f=100$ Hz (i.e., middle column in Figure 2) under geometric optics are shown in Figure 3. Comparing Figure 2 and 3, we can see that inside the diamond caustic as well as far from it in outer regions, the positions of de-amplified regions are well-traced even under the geometric optics approximation. As we cross the fold caustic (from inside to outside), two of the extrema points of the arrival time delay surface merge and disappear, which explains the sudden change in the $|F|$ maps under geometric optics approximation in Figure 3. On the other hand, considering the full wave-optics solution, we can see that there is still a non-zero contribution from regions close to the points where the two images merged, as the distortion of the arrival time delay surface is a continuous function of the source position. In addition, we can also appreciate the effect of macro-magnification, $\mu_{\rm m}$. An increase in $\mu_{\rm m}$ effectively brings the oscillations in $F(f)$ at lower frequencies.
@@
Example amplification curves for source positions marked by black solid stars in the top row, black solid stars in the middle row, and green solid stars in the middle row of Figure 2 are shown in the top, middle, and bottom rows of Figure 4, respectively. For the source outside the diamond caustic and close to the cusp point in the top row of Figure 2, we have the formation of two virtual images with the minima being highly amplified. And as we saw in the point mass lens case, interference of two images with very different amplification factors, although it leads to frequency-dependent oscillations, the amplitude of these oscillations is small, as seen in the blue curve in the top row of Figure 4. We also notice that solutions under geometric optics differ significantly compared to the full wave optics solution. This difference between geometric and wave optics solutions is even more prominent for the orange curve representing a source inside the diamond caustic, which has three images with high magnification factors in geometric optics. That said, in the top panel of Figure 4, we see that the position of crests and troughs along the frequency axis matches in wave and geometric optics.
@@
In the middle row of Figure 4, we show amplification factors for source positions such that they follow a de-amplified region around the diamond caustic along the major axis, whereas the bottom row represents sources that go perpendicular to the de-amplified region. The merger and disappearance of images can be seen in the $\tilde{F}(t)$ curves and the corresponding inset plots. Here, for a source inside the diamond caustic, we again see that the position of crests and troughs along the frequency axis matches in wave and geometric optics. However, once we cross the caustic, the geometric optics solution loses contribution from two images and differs significantly from the full wave optics solution.
@@ 6 Binary lens
where $q_{i}=M_{i}/M_{T}$ with $M_{T}=M_{1}+M_{2}$ represents the mass ratios of primary and secondary components and $2|\boldsymbol{L}|$ is the distance between the two lens components. Here, we have assumed that both point mass lenses are at equal distances from the origin. The corresponding arrival time delay surface is given as,
@@
where various quantities have their usual meaning. Without loss of generality, we can always choose a coordinate system such that both lenses lie on the x-axis, i.e., $\boldsymbol{L}=(L,0)$. With that, we can see that a binary lens can be described by a total of five parameters, $(M_{1z},M_{2z},L,y_{1},y_{2})$. Note that the Chang-Refsdal lens is also a limiting case of the binary lens (Schneider & Weiss, 1986; Dominik, 1999).
The interference patterns for a binary lens with equal mass ratio components and different $L$ values are shown in Figure 5. Example amplification factor curves for some randomly chosen source positions, marked by a solid stars, are shown in Figure 6. Again, as we have seen in the last section, amplification factor curves can differ significantly compared to geometric optics approximation close to caustics, where a number of virtual images changes (top row in Figure 6). As we move away from caustics, both the amplitude and phase of the oscillation start to converge towards geometric optics results (bottom row in Figure 6). For comparison, in the bottom panel, we also show the point mass amplification factor with $y=|\boldsymbol{y}|$. The interference patterns for a binary lens with a mass ratio  $(q_{1},q_{2})=(0.25,0.75)$ are shown in Figure 7. For a binary lens, again, the dimensionless frequency $(\nu=4{\rm G}M_{Tz}f/{\rm c^{3}})$, can be used to determine other pairs of total mass and frequency (assuming the same mass ratios) which would produce the above amplification maps.
@@ 7 Summary
In this work, we simulate and look at the patterns in the amplification maps for simple lens models, namely, the point mass lens, the Chang-Refsdal lens, and the binary lens, primarily focusing on GW signals detected with the LVK detector network. For the point mass lens, we have a point caustic at the centre of the source plane. The corresponding amplification factor calculated under geometric optics closely follows the full solution after the first peak (i.e., amplification region; see Bulashenko & Ubach 2022). However, once we go beyond the point mass lens, either by introducing external shear or by adding another point mass lens (i.e., binary lens), the corresponding caustics cover regions of different multiplicities in the source plane. As the source moves close to the caustics, the corresponding amplification factor shows significant differences compared to the amplification factor obtained under geometric optics approximation. Caustics also mark the region around which we have multiple virtual images with similar and high amplification, and can lead to significant de-amplification around ${\sim}100$ Hz for a lens mass of ${\sim}100{\rm M_{\odot}}-200{\rm M_{\odot}}$. If these significantly de-amplified regions in the amplification factor align with ringdown frequencies in the LVK band, we may end up with GW signals, which may be hard to interpret without the incorporation of lensing.
Another interesting aspect would be to search for unique features in the amplification factor profile for different lens models. Previous works have discussed the importance of unique features in breaking local degeneracies and distinguishing different lens models (e.g., Meena, 2024b). For example, a lens model with a finite core will lead to the formation of a central image, which introduces secondary modulations to the amplification factor (e.g., Tambalo etalet~al., 2023). Although less likely for GW signals with typical signal-to-noise values, such features may allow us to infer the presence of lensing and the lens properties without doing detailed parameter inferences.
@@ Acknowledgements
The author thanks the anonymous referee for useful comments. The author acknowledges the support of grant 2020750 from the United States-Israel Binational Science Foundation (BSF), grant 2109066 from the United States National Science Foundation (NSF), and the Ministry of Science & Technology, Israel. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.